OpenStudy (anonymous):

Find A and B (defined by x and y) (2^x*Square root3^y)/(2^(y-x)* 1/3^x) = 2^A * 3^B

OpenStudy (anonymous):

Is that 1 over 3^x?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

So it is actually 2 to the power[ (y-x) / 3^x], right?

OpenStudy (anonymous):

It's \[2^{x} * \frac{ 1 }{ 3^{x} }\]

OpenStudy (anonymous):

The second part, not the first....

OpenStudy (anonymous):

(2^ (x*Square root(3^y))) / (2^((y-x)* 1/3^x)) Is this right?

OpenStudy (anonymous):

\[\frac{ 2^{x}*\sqrt{3^{y}} }{ 2^{y-x} *\frac{ 1 }{ 3^{x} } } = 2^{A} * 3^{B}\] Thats the whole

OpenStudy (anonymous):

If so, since everything is to power 2 (top and bottom) you can just transfer the denominator to the numerator so 2 ^ (x * cubert (3y) - 3 power -x * (y-x)) and then just compare to find A and B

OpenStudy (anonymous):

That wont be right, seeing as you used 2 ^(x*...), or will it?

OpenStudy (anonymous):

Ah, I didn't see that you changed the equation......

OpenStudy (anonymous):

I had ^(x*cuberoot y) and you put (^x) * cuberoot y But still, the principle is same , bring everything up to the top eg 2^x and 2^(y-x) just become 2^y on top etc

OpenStudy (anonymous):

Ah, i start to see it now.

OpenStudy (anonymous):

Great:-)